OpenStudy (anonymous):

83=10^x+4 answer is -2.081?

OpenStudy (anonymous):

no its 1.89763.

OpenStudy (anonymous):

subtract 4 from both sides to get: 83-4 = 10^x+4-4 79 = 10^x writing this in logarithmic form base 10: x = log_10(79) punch that into the calculator to get an approximation.